if the line y= √3 x cuts the curve x 4 + ax 2y +bxy+cx+dy+6=0 at A,B,CandD then OA.OB.OC>OD(where O is the origin) is? ans. 96

if the line y=√3 x cuts the curve x4 + ax2y +bxy+cx+dy+6=0 at A,B,CandD then OA.OB.OC>OD(where O is the origin) is?


ans. 96

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

x4+√3ax3+√3bx2+cx+√3dx+6=0,,x1,x2,x3,x4,are the four roots,OA=sqrt(x1^2+y1^2)=sqrt(x1^2+3x1^2)=2x1,OB=2x2, OC=2x3, OD=2x4, OA.OB.OC.OD=16X1*X2*X3*X4=16*6=96